OpenStudy (anonymous):

Find the volume of the region bounded by z^2=x^2+y^2 and 2x=x^2+y^2. The first one is a paraboloid and the second is a cylinder.

OpenStudy (anonymous):

The only thing I can think of is \[\int\limits_{0}^{2}\int\limits_{-\pi/2}^{\pi/2}\int\limits_{r^2}^{rsintheta} r dr d \theta dz\]

OpenStudy (anonymous):

that gives me 4

OpenStudy (dan815):

is that supposed to be 2z=x^2+y^2

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its a cylinder that is unbounded by z

OpenStudy (anonymous):

its centered at x=1

OpenStudy (dan815):

oh ok i see

OpenStudy (dan815):